# Binary Autoencoder for Mechanistic Interpretability of Large Language Models

### v1

@@ 4.1 Evaluating Entropy Calculation from Binary Autoencoder
To evaluate the entropy estimation of BAE’s hidden binary activations ($h_{1}$), we build a synthetic random directional benchmark: (1) sample a $d$-dimensional $r$-rank orthonormal basis $M\in\mathbb{R}^{r\times d}$, (2) sample $r$ binary coefficients $c\in{0,1}^{r}$, and (3) generate an instance $cM$ by summing selected $M$ basis from $1$-elements in $c$. Repeat (2) and (3) $n$ times, we get a synthesis random directional dataset with $n$ samples. Intuitively, the entropy of the dataset is $r$ since the only randomness comes from the $r$ independent Bernoulli coefficients given the fixed basis. We train the BAE on a set of such datasets with various $r$ (details in Appendix A.1), and confirm whether the entropies calculated following the aforementioned method hit the corresponding $r$.
@@
Natural Low-entropy Tendency.
Moreover, interestingly, when the estimated entropy constraint is disabled (red in Fig. 2), the entropy gathers higher but near the diagonal, suggesting that BAE without entropy training objective can find a relatively low-entropy encoding $h_{1}$, even if no normalizations are conducted to minimize the encoding entropy. To get a closer observation, we plot the self-regression loss $\mathcal{L}_{r}$ and training-time entropy $\mathcal{L}_{e}$ on every training step, with $r=2$ (more cases in Appendix D.2), as shown in Fig. 3, with normal setting (left) or $\alpha_{e}=\alpha_{c}=0$ (right). In the normal setting (left), a stable and monotonous convergence to both minimal entropy and self-regression loss can be observed. However, if the entropy loss is disabled (right), the training dynamics become multiphase and non-monotonic. Specifically, Phase 1: the loss rapidly decreases while the hidden activations maintain relatively high entropy; Phase 2: entropy sharply drops to near zero, with the loss remaining low, even if no entropy and weight penalty are utilized; and Phase 3: both loss and entropy oscillate in a narrow range at the end of training. Such an observation suggests that: Gradient descent on simple regression loss finds relatively “simple” representations, after a long-term stagnation on loss value, supporting previous work on training dynamics (Tishby & Zaslavsky, 2015; Saxe et al., 2018; Huh et al., 2023; Nanda et al., 2023; Shah et al., 2020; Frankle & Carbin, 2019; Bartlett et al., 2020). Additionally, enabling the entropy penalty suppresses the harmful Phase 3 oscillations, confirming its effectiveness.
4.2 Tracing the Normal Sentence Modeling by Entropy
In this section, we track the LLM’s feed-forward dynamics by the entropy of hidden states. In detail, we sample $n=262144$ sentences from Pile (Gao et al., 2020), inputting them into Llama 3.2-1B [^3] Datasets and models in this paper are loaded from HuggingFace. (Grattafiori et al., 2024), then extract the hidden states of $2^{0,1,\cdots,10}$-th tokens from each layer. Then we train a BAE on each layer and position, with experiment details in Appendix A.2. From the trained BAE, we calculate the entropy as shown in Fig. 4, where we observe:
@@ 4.2 Tracing the Normal Sentence Modeling by Entropy
The entropy of hidden states from a specific layer increases with the number of prefix tokens and eventually saturates at a fixed value. This observation suggests that: if the hidden space of a specific layer is viewed as a channel (Elhage et al., 2021) through which token information is communicated, then the channel has a fixed bandwidth, limiting the amount of accommodated token information. That is, Transformers have implicit context windows on each layer, where the exceeded information is discarded or distorted. Unlike convolutional neural networks, we have no evidence to infer that such context windows are contiguous or even binary, so that it is possible to overwrite parts of but not the whole information from an old token by new information, causing distortation like the lost-in-the-middle problem (Hsieh et al., 2024; Liu et al., 2024a; He et al., 2024; An et al., 2024; Liu et al., 2025).
@@
As shown in Fig. 4, for a specific token position, hidden states from a deeper layer carry more information, and the entropy also saturates later on a deeper layer, suggesting that Transformer blocks sequentially inject contextualized information into hidden states. As a result, deeper layers are more likely to simulate broader context windows, which facilitates the processing of downstream tasks at higher levels of information aggregation and abstraction. In contrast, shallower layers, restricted by narrower context windows, tend to focus on local linguistic-level features and propagate this information to the later layer to combine into broader-scoped abstraction, which is consistent with intuition and previous empirical observations (Jawahar et al., 2019; Chen et al., 2023; Wang et al., 2023; Xiao et al., 2025). We discuss this point deeper in §6.
@@ 4.3 Tracing In-context Learning by Entropy
In this section, we track the LLM’s ICL (refer Appendix A.3 for introduction) inference dynamics via entropy, similar to §4.2. In detail, we sample $262144$ ICL input instances from SST-2 (Socher et al., 2013) with specific demonstration numbers, then extract the hidden states of the last token (the “:” in the “sentiment:”, where the answer to the query will be generated) of a specific layer, and then train a BAE on these hidden states, with prompt templates, settings and parameters detailed in Appendix A.3. We calculate the entropy as shown in Fig. 5, where: (1) similar to the normal language modeling, a deeper layer contains more information, which suggests that the implicit context windows hypothesis still stands for the ICL scenario. However, (2) contradicts the normal language modeling, hidden states with more demonstrations (globally longer sentence lengths and better accuracy, see Appendix A.3) counterintuitively contain lower entropy, which suggests an interesting conclusion: ICL is achieved through removal of information, where the information useless to the specified task may be suppressed by the given demonstrations. This contradicts the mainstream idea that ICL “learns new tasks or knowledge” (Pan et al., 2023; Li et al., 2024; Wang et al., 2025), and thus may offer a new perspective for interpreting the ICL inference processing as “deleting unrelated information from the query encoding (Cho et al., 2025) on the last token”.
@@ 5 Atomized Features Extraction by BAE
Given that our entropy constraint promotes sparsity and decorrelation of hidden activations $h_{1}$, the linear decomposition from $h_{0}$ to $h_{1}$ can disentangle atomized features in LLM hidden states (Elhage et al., 2022; Hänni et al., 2024). To evaluate such disentanglement, in this section, we assign human-interpretable semantics to each channel of $h_{1}$ and assess their consistency across inputs, using an improved automatic interpretation method under the LLM-as-a-judge framework (Gu et al., 2024).
5.1 Common Semantics-based Feature Interpretation and Evaluation
@@
We begin with a revisit to the current automatic interpretation and evaluation methods of features (Bills et al., 2023; Huben et al., 2024), where for a channel in $h_{1}$ corresponding to a feature (line vector) in $W_{\text{out}}$, (Step 1) given one input text, the activation magnitudes (e.g., the value of the specified channel in $h_{1}$ of SAE) on the channel are calculated on every token, then (Step 2) the tuple of sentence and activation magnitudes on all tokens is input in LLMs with a prompt (e.g., “Please predict the explanation of the feature given the following activations.”) for interpreting such feature into one phrase (e.g., “freedom-related terms”). Then, (Step 3) given a test input text, a simulation of the activation magnitudes based on the generated interpretation for each token is queried from the LLM, and the correlation coefficient of the simulated activation magnitudes and the SAE-calculated activation magnitudes is regarded as the interpretability score of this feature (refer to Bills et al. (2023) for details).
Notice that Step 2 of the above process relies on the LLM’s ability to accurately handle large amounts of numerical tokens to generate reliable explanatory phrases, and Step 3 requires the LLM to faithfully simulate activation magnitudes by numerical tokens, which places a high demand on the LLM’s mathematical reasoning capabilities and also output calibrations. However, current research indicates that LLMs exhibit weaker capabilities in numerical reasoning compared to linguistic tasks (Press et al., 2023; Schick et al., 2023; Ahn et al., 2024; Xu et al., 2025). Also, the output can be implicitly biased to some specific tokens (Zhao et al., 2021; Geng et al., 2024), making the aforementioned pipeline unreliable, requiring a revision to improve the robustness and credibility.
@@
Therefore, to avoid relying on LLMs to directly process numerical tokens, we propose ComSem as a new pipeline that leverages the LLMs’ strength in linguistic semantic recognition to interpret the extracted features. In detail (refer Appendix A.4 for the detailed pseudocode): for a specific channel (corresponding to a feature) in the $h_{1}$ of an autoencoder, given a set of test sentences, (Step 1) for all the tokens in the set, whose hidden states have significant activation magnitude on the specific channel (detailed in §5.2 for BAE), we collect them along with their sentences. (Step 2) We query the backend LLM to find the commonality of these tokens presented in their sentences, as the interpretation of the feature corresponding to the channel. (Step 3) On a hold-out set of such token-sentence tuples significantly activated on the channel, we query the LLM to judge whether these tokens can be interpreted by the generated interpretation from Step 2. The ratio of the “Yes” answer is calculated as the interpretability score of this feature. ComSem can avoid the concern in mathematical and numerical reasoning, and provides the possibility of applying simple output calibrations [^4] E.g., described in (Zhao et al., 2021; Fei et al., 2023). Such output calibration is not utilized in this work. on the true-or-false output. So, we utilize ComSem to evaluate the BAE, as described below.
5.2 Extracting, Interpreting and Evaluating BAE Features by ComSem
To utilize the ComSem pipeline on BAE, we first propose a rescaling method to identify the significant activations (utilized in Step 1 of ComSem) from BAE, since BAE’s binary hidden activations cannot directly reflect the activation magnitudes. Then, we evaluate the interpretability score of BAE-extracted features by ComSem to demonstrate that BAE is an effective feature extractor.
Estimate the Activation Magnitude from the Feature Burstiness.
@@ 5.2 Extracting, Interpreting and Evaluating BAE Features by ComSem
Since our aforementioned ComSem pipeline requires access to feature activation magnitude, which our BAE can not explicitly provide [^5] Note that, in typical SAE, the hidden activation (congener of our $h_{1}$) can directly serve as such magnitude., as shown in Fig. 10, we calculate the burstiness of each channel to convert the binary activations $h_{1}$ into the activation magnitudes. In detail, for each (indexed by $i$) instance $h_{1}^{(i)}$ in a hidden activation set $H_{1}=\{h_{1}^{(i)}\}_{i=1}^{n}$, we compare the distance of $h_{1}^{(i)}$ with a prior distribution [^6] Such $\bar{h_{1}}$ can either be an accompanying component saved during training, or mean value on $H_{1}$ set. $\bar{h_{1}}$ as $\beta^{(i)}=\log_{2}|h_{1}^{(i)}-\bar{h_{1}}|$, where the $\log_{2}$ is the element-wise logarithm. Such $\beta^{(i)}$ is the channel-wise activation magnitude of $h_{1}^{(i)}$, where a channel $j$ with larger $\beta^{(i)}_{j}$ suggests the $j$-th feature from $h_{1}^{(i)}$ is more bursty. A more bursty feature reduces more uncertainty to reconstruct the $h_{0}$ so that carries more information against the prior distribution $\bar{h_{1}}$, which makes it more representative of $h_{0}$. Then, for each instance, we simply select the indices with the top-$k$ greatest values in $\beta^{(i)}$ as the significant activation channels utilized in Step 1 of ComSem.
Figure 6: Feature activation frequency distribution of Layer 11 (more layers in Appendix D.3).
@@
(1) Global sparsity induced by low entropy. Notice that our entropy training objective shown in Eq. 4 is actually punishing the burstiness summed from all the input samples on all the channels among the minibatches, therefore, the calculated $\beta^{(i)}$ for each sample is numerically sparse in two directions, i.e., (i) for one instance $h_{1}^{(i)}$, most of elements in $\beta^{(i)}$ are suppressed to be sufficient small values, and (ii) for one feature channel $j$, most of $\beta^{(i)}_{j}$ are small values among $i$ among all instances, which is aligned to the sparsity assumption of mechanistic interpretability (refer to Appendix D.1 for visualization). (2) Consistent scale and clear interpretability. In typical SAEs, different channels of $h_{1}$ lie on varying scales, making some inherently high-magnitude channels appear consistently active (“dense features”, Appendix D.1 and Fig. 6(Stolfo et al., 2025; Rajamanoharan et al., 2024b) while others seem inactive (“dead features”). This issue is particularly pronounced when using top-$k$ selection based on absolute activation values to identify significant features, reducing the number and introducing confusion to activated features. Therefore, intuitively, as shown in §5.3, rescaling activations into an aligned distribution alleviates this problem in typical SAE. And in our pipeline, BAE simplifies the rescaling since each channel follows a Bernoulli distribution and requires only one statistic ($\bar{h_{1}}$), whereas the real-valued $h_{1}$ in typical SAEs is harder to characterize and rescale.
Table 1: Evaluation of BAE and baselines on 2 backend LLMs ($d^{\prime}/d=4$). Feature Activated: The number of features where sufficient $h_{0}$ instances (more than $8$, refer to Appendix A.4) with significant activation magnitude are observed on the corresponding channel. FI#: The number of feature channels with non-zero ComSem interpretability score. Score: The averaged ComSem score among all activated channels.
Feat.
Source
Model
(Appendix A.4)
Feature
Activated
$_{\text{4.1-mini}} ComSem$$_{\text{4.1}}$
@@
Llama 3.2-1B Layer 5 $d=2048$
ReLU SAE 1922 1103 0.216 1251 0.294
Top-k SAE 3531 2614 0.439 2763 0.445
Gated ReLU SAE 17 11 0.382 6 0.294
ReLU SAE Resc. 567 309 0.233 367 0.269
TransCoder 335 218 0.275 195 0.238
BAE (ours) 3012 2249 0.422 2449 0.439
Llama 3.2-1B Layer 9 $d=2048$
ReLU SAE 1914 1128 0.229 1274 0.254
Top-k SAE 3179 2389 0.450 2542 0.461
Gated ReLU SAE 100 85 0.638 80 0.596
ReLU SAE Resc. 3161 2028 0.258 2190 0.294
TransCoder 417 257 0.260 263 0.264
BAE (ours) 4675 3370 0.370 3624 0.394
Llama 3.2-1B Layer 11 $d=2048$
ReLU SAE 2065 1177 0.232 1380 0.260
Top-k SAE 3417 2540 0.440 2684 0.452
Gated ReLU SAE 1226 976 0.531 1026 0.557
ReLU SAE Resc. 744 435 0.259 482 0.278
TransCoder 1794 979 0.218 1090 0.239
BAE (ours) 5464 3882 0.360 4140 0.382
Llama 3.2-1B Layer 14 $d=2048$
ReLU SAE 2528 1423 0.195 1600 0.217
Top-k SAE 2702 1900 0.389 2004 0.418
Gated ReLU SAE 2948 2095 0.412 2250 0.435
ReLU SAE Resc. 3962 2361 0.252 2661 0.274
TransCoder 3401 1931 0.237 2166 0.267
BAE (ours) 6120 3963 0.324 3971 0.323
Llama 3.2-3B Layer 20 $d=3072$
ReLU SAE 1923 1183 0.289 1289 0.312
Top-k SAE 3234 2286 0.402 2508 0.425
Gated ReLU SAE 4628 3271 0.402 3580 0.437
ReLU SAE Resc. 2122 1222 0.250 1451 0.294
TransCoder 5508 3001 0.233 3424 0.257
BAE (ours) 9308 5956 0.308 6805 0.348

### v2

@@ 4.1 Evaluating Entropy Calculation from BAE
To evaluate the entropy estimation of BAE’s hidden binary activations ($h_{B}$), we build a synthetic random directional benchmark: (1) sample a $d$-dimensional $r$-rank orthonormal basis $M\in\mathbb{R}^{r\times d}$, (2) sample $r$ binary coefficients $c\in\{0,1\}^{r}$, and (3) generate an instance $cM$ by summing selected $M$ basis from $1$-elements in $c$. Repeat (2) and (3) $n$ times, we get a synthesis random directional dataset with $n$ samples. Intuitively, the entropy of the dataset is $r$ since the only randomness comes from the $r$ independent Bernoulli coefficients given the fixed basis. We train the BAE on a set of such datasets with various $r$ (details in Appendix A.1), and confirm whether the entropies calculated following the aforementioned method hit the corresponding $r$. Moreover, we propose a prototype to evaluate such entropy calculation on real LM hidden states, refer to Appendix B for details. Also, we analyse the training dynamics of BAE, and find a natural low-entropy trend of BAE hidden activations, refer to Appendix D.
@@
4.2 Tracing Language Modeling by Entropy
In this section, we track the LLM’s feed-forward dynamics by the entropy of hidden states. In detail, we sample $n=262144$ sentences from Pile (gao2020pile), inputting them into Llama 3.2-1B (grattafiori2024llama), then extract the hidden states of $2^{0,1,\cdots,10}$-th tokens from each layer. Then we train a BAE on each layer and position, with details in Appendix A.2. From the trained BAE, we calculate the entropy as shown in Fig. 3, where we observe:
@@ 4.2 Tracing Language Modeling by Entropy
The entropy of hidden states from a specific layer increases with the number of prefix tokens and eventually saturates at a fixed value. This observation suggests that: if the hidden space of a specific layer is viewed as a channel (elhage2021mathematical) through which token information is communicated, then the channel has a fixed bandwidth, limiting the amount of accommodated token information. That is, Transformers have implicit context windows on each layer, where the exceeded information is discarded or distorted. Unlike convolutional neural networks, we have no evidence to infer that such context windows are contiguous or even binary, so that it is possible to overwrite parts of but not the whole information from an old token by new information, causing distortation like the lost-in-the-middle problem (hsieh2024ruler; liu-etal-2024-lost; he-etal-2024-never; an2024make; liu2025comprehensivesurveylongcontext).
@@
As shown in Fig. 3, for a specific token position, hidden states from a deeper layer carry more information, and the entropy also saturates later on a deeper layer, suggesting that Transformer blocks sequentially inject contextualized information into hidden states. As a result, deeper layers are more likely to simulate broader context windows, which facilitates the processing of downstream tasks at higher levels of information aggregation and abstraction. In contrast, shallower layers, restricted by narrower context windows, tend to focus on local linguistic-level features and propagate this information to the later layer to combine into broader-scoped abstraction, which is consistent with intuition and previous empirical observations (jawahar-etal-2019-bert; chen2023which; wang2023label; xiao2025muddformer). We discuss this point deeper in §7.
@@ 4.3 Tracing In-context Learning by Entropy
In this section, we track the LLM’s ICL (refer Appendix A.3 for introduction) inference dynamics via entropy, similar to §4.2. In detail, we sample $262144$ ICL input instances from SST-2 (socher-etal-2013-recursive) with specific demonstration numbers, then extract the hidden states of the last token (the “:” in the “sentiment:”, where the answer to the query will be generated) of a specific layer, and then train a BAE on these hidden states, with prompt templates, settings and parameters detailed in Appendix A.3. We calculate the entropy as shown in Fig. 4, where: (1) similar to the normal language modeling, a deeper layer contains more information, which suggests that the implicit context windows hypothesis still stands for the ICL scenario. However, (2) contradicts the normal language modeling, hidden states with more demonstrations (globally longer sentence lengths and better accuracy, see Appendix A.3) counterintuitively contain lower entropy, which suggests an interesting conclusion: ICL is achieved through removal of information, where the information useless to the specified task may be suppressed by the given demonstrations. This contradicts the mainstream idea that ICL “learns new tasks or knowledge” (pan-etal-2023-context; li-etal-2024-language; wang2025can), and thus may offer a new perspective for interpreting the ICL inference processing as “deleting unrelated information from the query encoding (cho2025revisiting) on the last token”.
@@ 5 Atomized Features Extraction by BAE
Given that our entropy constraint promotes sparsity and decorrelation of $h_{B}$, the linear decomposition from $h_{0}$ to $h_{B}$ can disentangle atomized features in LLM hidden states (elhage2022toymodelsof; nni2024mathematical). To evaluate such disentanglement, in this section, we assign human-interpretable semantics to each channel of $h_{B}$ and assess their consistency across inputs, using an improved automatic interpretation method under the LLM-as-a-judge framework (gu2024survey). Also, we utilize BAE to extract steering vectors (hendel-etal-2023-context; sae_taskvectors) from ICL inputs and inject such steering vectors into 0-shot inputs to evaluate the effectiveness against SAE variants.
5.1 Common Semantics-based Feature Interpretation
@@
We begin with a revisiting to the current automatic interpretation and evaluation methods of features (bills2023language; huben2024sparse), where for a channel in $h_{B}$ corresponding to a feature (line vector) in $W_{\text{out}}$, (Step 1) given one input text, the activation magnitudes (e.g., the value of the specified channel in $h_{B}$ of SAE) on the channel are calculated on every token, then (Step 2) the tuple of sentence and activation magnitudes on all tokens is input in LLMs with a prompt (e.g., “Please predict the explanation of the feature given the following activations.”) for interpreting such feature into one phrase (e.g., “freedom-related terms”). Then, (Step 3) given a test input text, a simulation of the activation magnitudes based on the generated interpretation for each token is queried from the LLM, and the correlation coefficient of the simulated activation magnitudes and the SAE-calculated activation magnitudes is regarded as the interpretability score of this feature (refer to (bills2023language) for details).
Notice that Step 2 of the above process relies on the LLM’s ability to accurately handle large amounts of numerical tokens to generate reliable explanatory phrases, and Step 3 requires the LLM to faithfully simulate activation magnitudes by numerical tokens, which places a high demand on the LLM’s mathematical reasoning capabilities and also output calibrations. However, current research indicates that LLMs exhibit weaker capabilities in numerical reasoning compared to linguistic tasks (press2023measuring; schick2023toolformer; ahn2024large; xu2025can). Also, the output can be implicitly biased to some specific tokens (pmlr-v139-zhao21c; geng-etal-2024-survey), making the aforementioned pipeline unreliable, requiring a revision to improve the robustness and credibility.
@@
Therefore, to avoid relying on LLMs to directly process numerical tokens, we propose ComSem as a new pipeline that leverages the LLMs’ strength in linguistic semantic recognition to interpret the extracted features. In detail (refer Appendix A.4 for the detailed pseudocode): for a specific channel (corresponding to a feature) in the $h_{B}$ of an autoencoder, given a set of test sentences, (Step 1) for all the tokens in the set, whose hidden states have top-$k$ ($k$ is a hyperparameter of the method) significant activation magnitude on the specific channel over all channels (i.e., if the activation magnitude of this specific channel is in the top-$k$ when we sort all the activation magnitudes over all channels), we collect them along with their sentences. (Step 2) We query the backend LLM to find the commonality of these tokens presented in their sentences, as the interpretation of the feature corresponding to the channel. (Step 3) On a hold-out set of such token-sentence tuples significantly activated on the channel, we query the LLM to judge whether these tokens can be interpreted by the generated interpretation from Step 2. The ratio of the “Yes” answer is calculated as the interpretability score of this feature. ComSem can avoid the concern in mathematical and numerical reasoning, and provides the possibility of applying simple output calibrations [^2] E.g., described in (pmlr-v139-zhao21c; fei-etal-2023-mitigating). Such output calibration is not utilized in this work. on the true-or-false output. So, we utilize ComSem to evaluate the BAE, as described below.
5.2 Interpreting and Evaluating BAE Features
Estimate Activation Magnitude from Binary Activation.
@@ 5.2 Interpreting and Evaluating BAE Features
Since ComSem requires access to feature activation magnitude, which our BAE can not explicitly provide [^3] Note that, in typical SAE, the hidden activation (congener of our $h_{B}$) can directly serve as such magnitude., as shown in Fig. 10, we calculate the burstiness of each channel to convert the binary activations $h_{B}$ into the activation magnitudes. In detail, for each (indexed by $i$) instance $h_{B}^{(i)}$ in a hidden activation set $H_{B}=\{h_{B}^{(i)}\}_{i=1}^{n}$, we compare the distance of $h_{B}^{(i)}$ with a prior distribution [^4] Also named typical set in information theory. Such $\bar{h_{B}}$ can either be saved during training, or mean value on $H_{B}$ set. $\bar{h_{B}}\in[0,1]^{d^{\prime}}$ as $\beta^{(i)}=\log_{2}|h_{B}^{(i)}-\bar{h_{B}}|$, where the $\log_{2}$ is the element-wise logarithm. Such $\beta^{(i)}$ is the channel-wise activation magnitude of $h_{B}^{(i)}$, where a channel $j$ with larger $\beta^{(i)}_{j}$ suggests the $j$-th feature from $h_{B}^{(i)}$ is more bursty. A more bursty feature reduces more uncertainty to reconstruct the $h_{0}$ so that carries more information against the $\bar{h_{B}}$, which makes it more representative of $h_{0}$ (Appendix E for theoretical groundings). Then, we input such magitudes into the automatic feature interpretation and evaluation framework.
@@
(1) Global sparsity induced by low entropy. Notice that our entropy training objective shown in Eq. 4 is actually punishing the burstiness summed from all the input samples on all the channels among the minibatches, therefore, the calculated $\beta^{(i)}$ for each sample is numerically sparse in two directions, i.e., (i) for one instance $h_{B}^{(i)}$, most of elements in $\beta^{(i)}$ are suppressed to be sufficient small values, and (ii) for one feature channel $j$, most of $\beta^{(i)}_{j}$ are small values among $i$ among all instances, which is aligned to the sparsity assumption of mechanistic interpretability (refer to Appendix G.1 for visualization). (2) Consistent scale and clear interpretability. In typical SAEs, different channels of $h_{B}$ lie on varying scales, making some inherently high-magnitude channels appear consistently active (“dense features”, Appendix G.1 and Fig. 5(stolfo2025antipodal; rajamanoharan2024jumping) while others seem inactive (“dead features”). This issue is particularly pronounced when using top-$k$ selection based on absolute activation values to identify significant features, reducing the number and introducing confusion to activated features. Therefore, intuitively, as shown in §5.3, rescaling activations into an aligned distribution alleviates this problem in typical SAE. And in our pipeline, BAE simplifies the rescaling since each channel follows a Bernoulli distribution and requires only one statistic ($\bar{h_{B}}$), whereas the real-valued $h_{B}$ in typical SAEs is harder to characterize and rescale.
[…]
@@
Llama 3.2-1B Layer 5 $d=2048$ ReLU SAE $3{\times 10^{-3}}$ 0.08 1922 1103 0.216 1251 0.294
[…]
ReLU SAE Resc. $3{\times 10^{-3}}$ 0.08 567 309 0.233 367 0.269
[…]
TransCoder $2{\times 10^{-2}}$ 0.01 417 257 0.260 263 0.264
[…]
TransCoder $2{\times 10^{-2}}$ 0.01 1794 979 0.218 1090 0.239
[…]
TransCoder $0.15$ 0.01 5508 3001 0.233 3424 0.257
[…]